# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v3

@@ II-C1 Algorithm
II-C2 Problem
TABLE IV: Properties and methods of problem base class.
@@ 11 Prob = Prob@Problem(varargin);
All problems within MToP inherit from the problem base class <Problem.m>. The properties and methods of <Problem.m> are detailed in Table IV. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [8]. It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table LABEL:supp:tab:prob-mt of the supplementary files. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
II-C3 Performance Metric
@@
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is <MTOData>, which is generated during experimental execution. The function returns <result>, comprising properties such as <Metric>, <RowName>, <ColumnName>, <TableData>, <ConvergeData>, and <ParetoData>. The <Metric> property can take values of either ’Min’ or ’Max’, indicating whether a smaller or larger metric value is preferable. <RowName>, <ColumnName>, and <TableData> are utilized to present metric results in the GUI table. On the other hand, <ConvergeData> and <ParetoData> are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table V.
TABLE V: Metrics included in MToP. The table lists the metric name, optimization objective, task type, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. ST and MT represent single-task and multi-task, respectively
@@
IGD+ (MTS) MO MT Multitask score of IGD for all tasks
@@
III Performing Experiments with MToP
@@
Figure 6: Graphical user interface of MToP. The test module (a) is used for testing algorithms and problems. The graphical display is shown in the center section of the test module. The experiment module (b) is used for executing comparative experiments. The results of the experiment are displayed in the right section of the experiment module.
@@
III-A Testing Algorithms and Problems
@@
III-B Executing Comparative Experiments
@@
III-C Executing via Alternative Command Line
@@
IV Extending MToP
Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP. By adhering to the code pattern outlined in Section II-C, users can develop and implement their code files with ease.
IV-A Adding New Algorithms
@@ IV-B Adding New Problems
IV-C Adding New Performance Metrics
@@
IV-D Extending Optimization Types
There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking [18]. Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types. Users can refer to the existing code patterns in MToP to develop new optimization types. Specifically, for the sequential transfer optimization, it is possible to set the target task as the first task and the remaining tasks as source tasks for algorithms and problems. The metric calculation can be adjusted to evaluate the performance of the target task. For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
V Conclusion and Future Work
This paper introduces MToP, an open-source MATLAB platform designed for EMT. MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures. The current version of MToP contains over 40 MTEAs, more than 40 single-task EAs capable of handling MTO problems, over 150 benchmark MTO problems, and several real-world applications of EMT. This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
@@ V Conclusion and Future Work
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes [46, 4], multitask combinatorial optimization codes [5, 6] and more real-world applications [9, 47] in the future.
Acknowledgement
All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP.
The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO [36], for which the authors are also thankful.

### v4

@@ 2.3.1. Algorithm
Table 4. Multitask evolutionary algorithms included in MToP. The table lists the algorithm name, optimization objective, task type, framework, constraint handling, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. MT and MaT represent multi-task and many-task, respectively. MF and MP represent multifactorial and multi-population, respectively.
Algorithm Objective Task Framework Constraint Description
MFEA (Gupta et al., 2016) SO MT MF Multifactorial evolution-based GA for MTO
LDA-MFEA (Bali et al., 2017) SO MT MF Linearized domain adaptation in MFEA
MFDE&MFPSO (Feng et al., 2017) SO MT MF Multifactorial evolution-based DE and PSO
G-MFEA (Ding et al., 2019) SO MT MF Generalized MFEA with decision variable translation and shuffling
MFEA-GHS (Liang et al., 2019) SO MT MF Genetic Transform and hyper-rectangle search in MFEA
MFEA-DV (Yin et al., 2019) SO MT MF Enhanced MFEA with cross-task search direction
MFEA-II (Bali et al., 2020) SO MT MF MFEA with online transfer parameter estimation
SREMTO (Zheng et al., 2020) SO MT MF Self-regulated multitask framework for varying relatedness among tasks
TLTLA (Ma et al., 2020) SO MT MF MFEA with two-level inter- and intra-task transfer learning
MFEA-AKT (Zhou et al., 2021b) SO MT MF MFEA with adaptive knowledge transfer via multiple crossover operator
ASCMFDE (Tang et al., 2021a) SO MT MF Intertask transfer in aligned subspace for MFDE
AT-MFEA (Xue et al., 2022) SO MT MF Affine transformation-enhanced domain adaptation for MFEA
MFEA-DGD (Liu et al., 2023) SO MT MF MFEA based on diffusion gradient descent
MFEA-VC (Wang et al., 2024) SO MT MF MFEA with contrastive variational auto-encoder
IMEA (Hashimoto et al., 2018) SO MT MP Multi-population framework for MTO through island model
EMEA (Feng et al., 2019b) SO MT MP Explicit knowledge transfer for MTO via autoencoding
MFMP (Li et al., 2020) SO MT MP Multi-population-based adaptive DE for MTO
DEORA (Li et al., 2022c) SO MT MP Adaptive task selection for competitive MTO
MKTDE (Li et al., 2022b) SO MT MP Meta-knowledge transfer-based multitask DE
MTES (Bai et al., 2022) SO MT MP Multitask OpenAI-ES via gradient-free evolution multitasking
BLKT-DE (Jiang et al., 2023b) SO MT MP Multitask DE with block-level knowledge transfer
MTSRA (Li et al., 2023a) SO MT MP Improved adaptive DE with competitive task selection
MTDE-ADKT (Zhang et al., 2024a) SO MT MP Multitask DE with adaptive dual knowledge transfer
CEDA (Zhang et al., 2024b) SO MT MF/MP Constrained multitasking via co-evolution and domain adaptation
CMO-LKT (Ban et al., 2025) SO MT MP Local knowledge transfer-based EA for constrained MTO
MTEA-AD (Wang et al., 2022a) SO MT/MaT MP MTEA with adaptive knowledge transfer via anomaly detection
MTES-KG (Li et al., 2024c) SO MT/MaT MP Multitask ES with knowledge-guided external sampling
MTEA-HKTS (Zhao et al., 2024) SO MT/MaT MP MTEA for solving the problem of transfer targets
BoKT (Jiang et al., 2023a) SO MaT MP Bi-objective knowledge transfer framework for many-task optimization
TRADE (Wu et al., 2023) SO MaT MP Transferable adaptive parameter DE for many-task optimization
TNG-NES (Li et al., 2024b) SO MaT MP Transfer task-averaged natural gradient for efficient many-task optimization
MTEA-SaO (Li et al., 2023b) SO/MO MT/MaT MP Adaptive solver multitask framework with implicit knowledge transfer
KR-MTEA (Cui et al., 2023) SO/MO MT/MaT MP Adaptive multi-task EA based on knowledge reuse
SBO (Liaw and Ting, 2019) SO/MO MaT MP Symbiosis in biocoenosis framework for many-task optimization
MaTDE (Chen et al., 2020) SO/MO MaT MP Many-task DE with adaptive archive-based knowledge transfer
EMaTO-MKT (Liang et al., 2022c) SO/MO MaT MP Multi-source knowledge transfer via local distribution estimation
MO-MFEA (Gupta et al., 2017) MO MT MF Multiobjective MFEA with non-dominated sort and crowding distance
MO-MFEA-II (Bali et al., 2021) MO MT MF Cognizant Multitasking for parameter estimation in MO-MFEA
EMT-PD (Liang et al., 2022b) MO MT MF Two-stage adaptive knowledge transfer based on population distribution
EMT-GS (Liang et al., 2023) MO MT MF Generative adversarial networks for knowledge transfer
MM-DE (Chen et al., 2018) MO MT MP Fast memetic DE for multiobjective MTO
AMT-NSGA-II (Da et al., 2019) MO MT MP Curbing negative influences online for seamless transfer
EMT-ET (Lin et al., 2021) MO MT MP Effective knowledge transfer approach via non-dominated sort
MOMFEA-SADE (Liang et al., 2022a) MO MT MP Subspace alignment and adaptive Differential Evolution
MTEA-D-DN (Wang et al., 2023) MO MT MP Neighborhood as a bridge for decomposition-based knowledge transfer
MTEA-D-TSD  (Li et al., 2024a) MO MT MP Transfer search directions among decomposed subtasks
MTDE-MKTA (Li and Gong, 2025) MO MT MP Multitask DE with multiple knowledge types and transfer adaptation
RVC-MTEA (Li et al., 2024g) MO MT MP Reference vector contribution based MTEA for competitive multiobjective MTO
MTEA-DCK (Li et al., 2025) MO MT MP Multiobjective MTEA with diversity- and convergen-based knowledge transfer
Table 5. Single-task evolutionary algorithms for multitask optimization included in MToP. The table lists the algorithm name, optimization objective, task type, category, constraint handling, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. GA denotes genetic algorithm, DE denotes differential evolution, SI denotes swarm intelligence, ES denotes evolution strategy, CH denotes constraint handling, ND denotes non-dominated sorting, ID denotes indicator-based, DC denotes decomposition, and RV denotes reference vector.
Algorithms Objective Task Category Constraint Description
GA (Zbigniew, 1996) SO ST/MT GA Genetic algorithm with SBX crossover and polynomial mutation
Global-GA (Bull and Liu, 2024) SO ST/MT GA GA with global crossover to replace cooperative coevolution
DE (Storn, 1996) SO ST/MT DE Differential evolution algorithm with DE/rand/1/bin operator
jDE (Brest et al., 2006) SO ST/MT DE Self-adapting control parameters in DE
JADE (Zhang and Sanderson, 2009) SO ST/MT DE Adaptive DE with fast convergence performance
rank-DE (Gong and Cai, 2013) SO ST/MT DE Ranking-based mutation operators for DE
SHADE (Tanabe and Fukunaga, 2013) SO ST/MT DE Success-history based parameter adaptation for DE
L-SHADE (Tanabe and Fukunaga, 2014) SO ST/MT DE Linear population size reduction for SHADE
LSHADE44 (Poláková, 2017) SO ST/MT DE L-SHADE with competing strategies applied to constrained optimization
CAL-SHADE (Zamuda, 2017) SO ST/MT DE Adaptive constraint handling technique for L-SHADE
jSO (Brest et al., 2017) SO ST/MT DE Champion of CEC 2017 numerical optimization competition via improved DE
NL-SHADE-RSP (Stanovov et al., 2021) SO ST/MT DE Champion of CEC 2021 numerical optimization competition via improved DE
PSO (Kennedy and Eberhart, 1995) SO ST/MT SI Particle swarm optimization with global and particle best update
CSO (Cheng and Jin, 2015) SO ST/MT SI Competitive swarm optimizer for large-scale optimization
MPA (Faramarzi et al., 2020a) SO ST/MT SI Marine predators algorithm for numerical optimization
EO (Faramarzi et al., 2020b) SO ST/MT SI Equilibrium optimizer for numerical optimization
AO (Abualigah et al., 2021) SO ST/MT SI Aquila optimizer for numerical optimization
KLDE&KLPSO (Jiang et al., 2023c) SO ST/MT DE/SI Knowledge learning-based DE and PSO for numerical optimization
CMA-ES (Hansen and Ostermeier, 2001) SO ST/MT ES Evolution strategy with derandomized covariance matrix adaptation
IPOP-CMA-ES (Auger and Hansen, 2005) SO ST/MT ES Restart CMA-ES with increasing population size
sep-CMA-ES (Ros and Hansen, 2008) SO ST/MT ES Separable CMA-ES for large-scale optimization
xNES-as (Schaul, 2012) SO ST/MT ES Natural ES with adaptation sampling
R1-NES (Sun et al., 2013) SO ST/MT ES A Linear Time natural ES for non-separable functions
xNES&SNES (Wierstra et al., 2014) SO ST/MT ES ES with adaptive natural gradients to update distribution
OpenAI-ES (Salimans et al., 2017) SO ST/MT ES Parallelized ES with standard normal distribution gradients
DES (Arabas and Jagodziński, 2020) SO ST/MT ES Matrix-free covariance matrix adaptation ES
MTV-DE (Mezura-Montes et al., 2007) SO ST/MT CH Multiple trial vectors in DE for handling constraints
ECHT (Mallipeddi and Suganthan, 2010b) SO ST/MT CH Ensemble of constraint handling techniques
FROFI (Wang et al., 2016) SO ST/MT CH Incorporating objective function information into the feasibility rule
C2oDE (Wang et al., 2019) SO ST/MT CH Composite DE for constrained optimization
CORCO (Wang et al., 2020) SO ST/MT CH Utilizing the correlation between constraints and objective function
DeCODE (Wang et al., 2021) SO ST/MT CH Decomposition-based multiobjective approach for constrained optimization
VMCH (Wu et al., 2022) SO ST/MT CH Voting-mechanism-based ensemble of constraint handling techniques
CEDE-DRL (Hu et al., 2023) SO ST/MT CH Deep reinforcement learning assisted co-evolutionary DE
CCEF-ECHT (Li et al., 2024d) SO ST/MT CH Competitive and cooperative ensemble of constraint handling techniques
SPEA2 (Zitzler et al., 2001) MO ST/MT ND Improving the strength Pareto approach for multiobjective optimization
NSGA-II (Deb et al., 2002) MO ST/MT ND Multiobjective GA with non-dominated sort and crowding distance
NSGA-III (Deb and Jain, 2014) MO ST/MT ND Reference-point-based nondominated sorting for multiobjective optimization
SMS-EMOA (Beume et al., 2007) MO ST/MT ID Hypervolume-based selection for multiobjective optimization
MO-CMA-ES (Igel et al., 2007) MO ST/MT ES Covariance matrix adaptation for multiobjective optimization
MOEA/D (Zhang and Li, 2007) MO ST/MT DC Decomposition-based multiobjective evolutionary algorithm
MOEA/D-DE (Li and Zhang, 2009) MO ST/MT DC MOEA/D for complex Pareto sets in multiobjective optimization
RVEA (Cheng et al., 2016) MO ST/MT RV Reference vector-guided EA for many-objective optimization
AR-MOEA (Tian et al., 2018) MO ST/MT ID Adaptive reference points for MOEA/D in multiobjective optimization
MSEA (Tian et al., 2021a) MO ST/MT ND Multistage EA for better diversity preservation in multiobjective optimization
CCMO (Tian et al., 2021b) MO ST/MT CH Coevolutionary framework for constrained multiobjective optimization
LMOCSO (Tian et al., 2020) MO ST/MT SI Efficient large-scale multiobjective optimization based on CSO
CMOCSO (Ming et al., 2023) MO ST/MT SI Competitive and cooperative swarm optimizer for multiobjective optimization
2.3.2. Problem
All problems within MToP inherit from the problem base class <Problem.m>. The properties and methods of <Problem.m> are detailed in Table 7. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach (Gupta et al., 2016). It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table 6. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
Table 6. Benchmark and real-world optimization problems included in MToP. The table lists the problem name, case number, task type, dimension, optimization objective, task type, special case, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. ST, MT and MaT represent single-task, multi-task and many-task, respectively.
Problem Case Task Dimension Objective Task Special Description
CEC17-MTSO (Da et al., 2017) 9 2 50 SO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MaTSO (Feng et al., 2019a) 6 2-50 50 SO MaT Single-objective many-task optimization problems
WCCI20-MTSO (Feng et al., 2020) 10 2 50 SO MT Complex single-objective MTO problems
WCCI20-MaTSO (Feng et al., 2020) 10 2-50 50 SO MaT Complex single-objective many-task optimization problems
CMT (Li et al., 2022a) 9 2 Any SO MT Constrained Large/partial/no feasible regions interaction and different inter-task similarity
C2TOP&C4TOP (Li et al., 2022c) 51 2/4 50 SO MT Competitive Different optimal objective values among competitive tasks
C-CPLX (Li et al., 2023a) 20 2 50 SO MT Competitive Complex competitive MTO problems
LSMaTSO (Li et al., 2024b) 5 2-50 300 SO MaT Large-scale many-task optimization problems
CEC17-MTMO (Yuan et al., 2017) 9 2 50 MO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MTMO (Feng et al., 2019a) 10 2 50 MO MT Complex multi-objective MTO problems
CEC19-MaTMO (Feng et al., 2019a) 10 2-50 50 MO MaT Multi-objective many-task optimization problems
WCCI20-MaTMO (Feng et al., 2020) 10 2-50 50 MO MaT Complex multi-objective many-task optimization problems
CEC21-MTMO (Feng et al., 2021b) 10 2 50 MO MT More complex multi-objective multi optimization problems
CMOMT (Li et al., 2024g) 12 2/3 30-60 MO MT Competitive multiobjective MTO problems
Synthetic Functions 9 1 Any SO ST Ackley, Elliptic, Griewank, Rastrigin, Rosenbrock Schwefel, Sphere, Weierstrass
CEC06-CSO (Liang et al., 2006) 24 1 2-24 SO ST Constrained CEC 2006 competition on constrained single-objective optimization
CEC10-CSO (Mallipeddi and Suganthan, 2010a) 18 1 10/30 SO ST Constrained CEC 2010 competition on constrained single-objective optimization
CEC17-CSO (Wu et al., 2017) 28 1 2-100 SO ST Constrained CEC 2017 competition on constrained single-objective optimization
CEC17-SO (Awad et al., 2017) 29 1 2-100 SO ST CEC 2017 competition on bound constrained single-objective optimization
CEC20-RWCO (Kumar et al., 2020) 57 1 2-158 SO ST Constrained CEC 2020 competition on real-world constrained single-objective optimization
CEC22-SO (Kumar et al., 2022) 12 1 10/20 SO ST CEC 2022 competition on bound constrained single-objective optimization
PEPVM (Li et al., 2023b) 1 3 7 SO MT Parameter extraction of photovoltaic models as MTO problem
SCP (Li et al., 2022c) 1 Any 60-120 SO MT Competitive Sensor coverage problem as competitive MTO problem
OPF (Li et al., 2024c) 5 2 24/33 SO MT Constrained Optimal power flow as constrained MTO problem
PKACP (Jiang et al., 2023a) 1 Any Any SO MaT Planar kinematic arm control problem as many-task optimization problem
MO-SCP (Li et al., 2023b) 2 Any 60-120 MO MT Competitive Multiobjective sensor coverage problem as competitive multiobjective MTO problem
SOPM (Li and Gong, 2025) 2 3 30 MO MT Constrained Synchronous optimal pulse-width modulation as constrained multiobjective MTO problem
MO-OPF (Li and Gong, 2025) 4 2 24/33 MO MT Constrained Multiobjective optimal power flow as constrained multiobjective MTO problem
OPF-CMTMO (Li et al., 2024g) 4 2 11/24 MO MT Competitive Competitive multitask optimal power flow with thermal and wind-solar power
Table 7. Properties and methods of problem base class.
@@ 11 Prob = Prob@Problem(varargin);
2.3.3. Performance Metric
@@
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is <MTOData>, which is generated during experimental execution. The function returns <result>, comprising properties such as <Metric>, <RowName>, <ColumnName>, <TableData>, <ConvergeData>, and <ParetoData>. The <Metric> property can take values of either ’Min’ or ’Max’, indicating whether a smaller or larger metric value is preferable. <RowName>, <ColumnName>, and <TableData> are utilized to present metric results in the GUI table. On the other hand, <ConvergeData> and <ParetoData> are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table 8.
Table 8. Metrics included in MToP. The table lists the metric name, optimization objective, task type, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. ST and MT represent single-task and multi-task, respectively
@@
IGD (MTS) MO MT Multitask score of IGD for all tasks
@@
3. Performing Experiments with MToP
@@
Figure 6. Graphical user interface of MToP. The test module (a) is used for testing algorithms and problems. The graphical display is shown in the center section of the test module. The experiment module (b) is used for executing comparative experiments. The results of the experiment are displayed in the right section of the experiment module.
@@
3.1. Testing Algorithms and Problems
@@
3.2. Executing Comparative Experiments
@@
3.3. Executing via Alternative Command Line
@@
4. Extending MToP
Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP. By adhering to the code pattern outlined in Section 2.3, users can develop and implement their code files with ease.
4.1. Adding New Algorithms
@@ 4.2. Adding New Problems
4.3. Adding New Performance Metrics
@@
4.4. Extending Optimization Types
There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking (Tan et al., 2021). Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types. Users can refer to the existing code patterns in MToP to develop new optimization types. Specifically, for the sequential transfer optimization, it is possible to set the target task as the first task and the remaining tasks as source tasks for algorithms and problems. The metric calculation can be adjusted to evaluate the performance of the target task. For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
5. Conclusion and Future Work
This paper introduces MToP, an open-source MATLAB platform designed for EMT. MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures. The current version of MToP contains over 50 MTEAs, more than 50 single-task EAs capable of handling MTO problems, over 200 benchmark MTO problems, and several real-world applications of EMT. This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
@@ 5. Conclusion and Future Work
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng et al., 2016; Deb and Jain, 2014), multitask combinatorial optimization codes (Feng et al., 2021c, a) and more real-world applications (Gupta et al., 2022; Huang et al., 2023) in the future.
Acknowledgements.
This work was partly supported by the National Natural Science Foundation of China under Grant No. 62076225 and 62403477, the Open Research Fund of the State Key Lab of Digital Manufacturing Equipment & Technology under Grant No. IMETKF2024020, and the Major Project of Hubei Province Science and Technology under Grant No. 2023BCA006. All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP. The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian et al., 2017), for which the authors are also thankful.